if a security currently worth 9200 will be worth 11,193.21 5 years in the future what is the implied interest rate the investor will earn on the security assuming that no additional deposits or withdrawls are made.
Answer #1

The equation connecting Present value, PV and Future Value, FV of a security, assuming that no additional deposits or withdrawls are made, is given by

PV = FV / (1+ R)N

where R = interest rate and N is the period.

Hence, 9,200 = 11,193.21 / (1 + R)5

Hence, R = (11,193.21 / 9,200)1/5 - 1 = 4.00%.

Hence, the implied interest rate the investor will earn on the security = R = 4.00% per annum.